OpenStudy (anonymous):

x squared + 11x +10= 0

OpenStudy (anonymous):

do you know how to factor?

OpenStudy (anonymous):

right nope can't say i do...but it does need to be factored

OpenStudy (anonymous):

ok. consider this quadradic equation \[x^2+5x+6=0\] this factors to \[(x+3)(x+2)=0\] since we know that if 2 factors multiplied to =0 then one of them have to be zero so we do the following \[x+3=0\] or \[x+2=0\] does any of this help you

OpenStudy (anonymous):

can't say it does

OpenStudy (anonymous):

well, do a little work here. im not going to work it for you. notice that the 3 and 2 multiplied are 6 and the sum of 3 and 2 are 5

OpenStudy (anonymous):

please elaborate

OpenStudy (anonymous):

in my two binomials above i have (x+3)(x+2)=0 2 times 3 is 6 2 plus 3 is 5

OpenStudy (anonymous):

now look at my original equation x^2+5x+6. see the relationship?

OpenStudy (anonymous):

you know just call me retarded i still don't understand

OpenStudy (anonymous):

i dont think youre retarted. you just must not pay attention in class...

OpenStudy (anonymous):

yeah you can say that it's kinda hard to pay attention with a bunch of jar heads with you

OpenStudy (anonymous):

jar head?

OpenStudy (anonymous):

marines

OpenStudy (anonymous):

could you please help?

OpenStudy (whpalmer4):

Okay, here's the drill. \[x^2+11x+10=0\]You want to factor the left hand side. Take the coefficient (the number in front of the variable) of the first term, which is 1 (if there's no number, it's 1), and multiply it by the constant term, which is 10: 1*10 = 10. Now, you need to find two numbers that multiplied give you 10, and added, give you 11. What are those two numbers?

OpenStudy (anonymous):

dumb it up a little please

OpenStudy (whpalmer4):

two numbers that add up to 11. give me a pair.

OpenStudy (anonymous):

so i would do a headphone problem right?

OpenStudy (whpalmer4):

I don't know what a headphone problem is. Just tell me a pair of numbers that add to 11. 3+8 = 11. However, 3*8 = 24, not 10. Pick a different pair of numbers that add to 11.

OpenStudy (anonymous):

 headphones?